the ratio of poseurs to outright frauds was 14 to 3. if they totaled 2244, how many were poseurs
Answer by richwmiller(17219)

14+3=17 parts
14*2244/17=1848 poseurs
3*2244/17=396 frauds
1848/396=14/3
1848 poseurs
frauds 396